MEDICINE When doctors prescribe medication, they have to consider the rate at which the body filters a drug from the bloodstream. Suppose it takes the human body 6 days to filter out half of the Flu-B-Gone vaccine. The amount of Flu-B-Gone vaccine remaining in the bloodstream x days after an injection is given by the equation y = 𝑦0(0.5)^(𝑥/6), where 𝑦0 is the initial amount. Suppose a doctor injects a patient with 20 μg (micrograms) of Flu-B-Gone.
a. How much of the vaccine will remain after 1 day? Round your answer to the nearest tenth.
b. How much of the vaccine will remain after 12 days? Round your answer to the nearest tenth.
c. After how many days will the amount of vaccine be less than 1 μg?

Answer by Boreal(15235)

a. y=20*0.5^(1/6)=17.8 microgm
b. y=20*0.5^2=5.00 microgm
c. 1=20*0.5^(x/6)
0.05=0.5^(x/6)
ln both sides
ln 0.05=(x/6) ln 0.5
-3=(x/6)*(-0.693)
4.33=x/6
x=25.98 or 26.0 days